# publications.bib

### v3

@@ 2.2.2. Igusa Cusp Labels
In order to understand the boundary components $\partial\mathrm{Ig}^{b,*}$ and $\partial\mathrm{Ig}^{b,\mathrm{tor}}$, we will recall the notion of Igusa cusp labels, as in [San, Definition 3.2.19] (which we have slightly modified to match the definition of cusp labels previously introduced). We let $\mathbb{X}_{b}:=\mathbb{X}$ be the completely slope divisible $p$-divisible group attached to $b$ defined above. We reintroduce $b$ in the notation to emphasise that all constructions here depend on $b$. Finally, observe that, since from the moduli problem the polarization on $\mathcal{A}_{x}$ is prime-to-$p$, the $p$-divisible group $\mathbb{X}_{b}=\mathcal{A}_{x}[p^{\infty}]$ is principally polarized.
@@
    $Z_{b}$ is an $\mathcal{O}_{B}$stable filtration of $\mathbb{X}_{b}$ by $p$-divisible subgroups of the form
@@
    where $\mathrm{Gr}_{-2}^{Z_{b}}=Z_{b,-2}$ is multiplicative, and $\mathrm{Gr}_{0}^{Z_{b}}=\mathbb{X}_{b}/Z_{b,-1}$ is étale, and $Z_{b,-1}$, $Z_{b,-2}$ are Cartier dual to each other under the principal polarization on $\mathbb{X}_{b}$.
@@
    $\delta_{b}$ is an $\mathcal{O}_{B}$-linear isomorphism
@@
    $Z^{p}$ is an $\mathcal{O}_{B}$stable split, symplectic and admissible filtration
@@
There is an action of $J_{b}(\mathbb{Q}_{p})\times\mathbf{G}(\mathbb{A}^{p}_{f})$ on Igusa cusp labels. If $K\subset J_{b}(\mathbb{Q}_{p})\times\mathbf{G}(\mathbb{A}^{p}_{f})$ is a compact open subgroup then an Igusa cusp label at level $K$ is a $K$-orbit of Igusa cusp labels. For a general closed subgroup $H\subset J_{b}(\mathbb{Q}_{p})\times\mathbf{G}(\mathbb{A}^{p}_{f})$, an Igusa cusp label at level $H$ is a compatible family of Igusa cusp labels at level $K$ for all $K\supset H$.
@@ 2.2.3. Boundary components
We can decompose the boundary $\partial\mathrm{Ig}^{b,\mathrm{tor}}$ according to Igusa cusp labels of prime-to-$p$ level $K^{p}(N)$, in the following way. For every positive integer $m$, there is a level $p^{m}$-Igusa variety $\mathrm{Ig}^{b}_{m}$, defined as in [CS1, Definition 4.3.6], and we let $\Gamma_{m,b}$-denote the Galois group of the finite étale cover $\mathrm{Ig}^{b}_{m}\rightarrow\mathscr{C}_{\mathbb{X}}$. We also have a toroidal extension of $\mathrm{Ig}^{b}_{m}$ to a level $p^{m}$-Igusa variety $\mathrm{Ig}^{b,\mathrm{tor}}_{m}$, as defined in [CS2, Definition 3.2.5]. We let $\Gamma_{b}(p^{m}):=\mathrm{ker}(\mathrm{Aut}(\mathbb{X})\rightarrow\Gamma_{m,b})$, and note that if we let $K=\Gamma_{b}(p^{m})K^{p}(N)$ then such Igusa cusp labels at level $K$ have the same data as triples $(Z_{m,b},Z,\Phi)$, where $Z=(Z,\Phi)$ is a cusp label as level $K(N)$, and $Z_{m,b}$ is an $\mathcal{O}_{B}$-filtration on $\mathbb{X}_{b}[p^{m}]$ together with an isomorphism $X/p^{m}\simeq\mathrm{Gr}_{0}^{Z_{m,b}}$. In particular, we can consider the locally closed boundary stratum $\mathscr{C}_{\mathbb{X},Z}$, and the $p^{m}$-Igusa variety $\mathrm{Ig}^{b,\mathrm{tor}}_{m,Z}$ which is the preimage of $\mathscr{C}_{\mathbb{X},Z}$. Moreover, from [San, Theorem 3.2.22], we see that we have a decomposition
$\mathrm{Ig}^{b,\mathrm{tor}}_{m,Z}=\coprod_{\tilde{Z}_{m}}\mathrm{Ig}^{b,\mathrm{tor}}_{m,\tilde{Z}_{m}},$
@@
where $\tilde{Z}_{m}$ denotes Igusa cusp labels of level $\Gamma_{b}(p^{m})K^{p}(N)$ lying over $Z$.
Now, let $\tilde{Z}$ denote an Igusa cusp label of level $K^{p}(N)$. This is by definition a compatible system $\{\tilde{Z}_{m}\}$ of Igusa cusp labels for $\Gamma_{b}(p^{m})K^{p}(N)$, for all positive integers $m$. We can thus define
$\mathrm{Ig}^{b,\mathrm{tor}}_{\tilde{Z}}=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\leftarrow\\ m\end{subarray}}\mathrm{Ig}^{b,\mathrm{tor}}_{m,\tilde{Z}_{m}}.$
For later use, we will want to have a moduli description of points in $\mathrm{Ig}^{b,\mathrm{tor}}_{\tilde{Z}}$. We first recall some facts about degenerations of abelian schemes, from [Lan, §3.4] and [CS2, §2.5.1]. Let $C^{\prime}$ be a complete algebraically closed nonarchimedean field with ring of integers $\mathcal{O}_{C^{\prime}}$. Consider a polarized abelian variety $(A,\lambda)$ over $C^{\prime}$ with $\mathcal{O}_{B}$-structure, and a degeneration $\mathcal{A}$ of $A$ over $\mathcal{O}_{C^{\prime}}$. Then this uniquely determines a short exact sequence
@@
where $T$ is a torus, $\mathcal{B}$ is an abelian scheme over $\mathcal{O}_{C^{\prime}}$, and $\mathcal{G}$ is the Raynaud extension. Let $X=\mathbb{X}_{*}(T)$, which is a free abelian group over $\mathcal{O}_{C}$. The lattice $X$ has an action of $\mathcal{O}_{B}$, and so does $\mathcal{B}$. Then, $\mathcal{G}$ determines, and is uniquely determined by, an $\mathcal{O}_{B}$-linear map $c:X\rightarrow\mathcal{B}^{\vee}$. Similarly, we can consider a degeneration $\mathcal{A}^{\vee}$ over $\mathcal{O}_{C^{\prime}}$ of the dual $A^{\vee}/C^{\prime}$, which gives us a short exact sequence
@@
and if we similarly let $Y=\mathbb{X}_{*}(T^{\vee})=\mathbb{X}^{*}(T)$, the extension $\mathcal{G}$ determines, and is uniquely determined by an $\mathcal{O}_{B}$-linear map $c^{\vee}:Y\rightarrow\mathcal{B}$.
Moreover, the polarization $\mathcal{G}\rightarrow\mathcal{G}^{\vee}$ determines and is uniquely determined by the data of the polarization on $\mathcal{B}$, and an injective $\mathcal{O}_{B}$-linear map $\phi:Y\rightarrow X$.
Given $\tilde{Z}$, an Igusa cusp label of level $K^{p}(N)$, we want to understand the $(C,\mathcal{O}_{C})$-valued points of $\mathfrak{Ig}^{b,\mathrm{tor}}$ which specialize to points in $\mathrm{Ig}^{b,\mathrm{tor}}_{\tilde{Z}}$. In particular, note that the data of the Igusa cusp label means we have fixed an $\mathcal{O}_{B}$-stable filtration
$0=Z_{b,-3}\subset Z_{b,-2}\subset Z_{b,-1}\subset\mathbb{X}_{b},$
as well as a $\mathcal{O}_{B}$linear splitting $\delta_{b}$ of this filtration. From the proof of [San, Theorem 4.3.10], we see that $(C^{\prime},\mathcal{O}_{C^{\prime}})$-valued points are hence given by the following data:

### v5

@@ 2.2.3. Igusa Cusp Labels
In order to understand the boundary components $\partial\mathrm{Ig}^{b,*}$ and $\partial\mathrm{Ig}^{b,\mathrm{tor}}$, we will recall the notion of Igusa cusp labels, as in [54, Definition 3.2.19] (which we have slightly modified to match the definition of cusp labels previously introduced, and which is similar to [8, Definition 3.3.10]). We let $\mathbb{X}_{b}:=\mathbb{X}$ be the completely slope divisible $p$-divisible group attached to $b$ defined above. We reintroduce $b$ in the notation to emphasise that all constructions here depend on $b$. Finally, observe that, since from the moduli problem the polarization on $\mathcal{A}_{x}$ is prime-to-$p$, the $p$-divisible group $\mathbb{X}_{b}=\mathcal{A}_{x}[p^{\infty}]$ is principally polarized.
@@
    $Z_{b}$ is an $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$ -stable filtration of $\mathbb{X}_{b}$ by $p$-divisible subgroups of the form
@@
    with quotients which are also $p$-divisible groups, where $\mathrm{Gr}_{-2}^{Z_{b}}=Z_{b,-2}$ is multiplicative, and $\mathrm{Gr}_{0}^{Z_{b}}=\mathbb{X}_{b}/Z_{b,-1}$ is étale, and $Z_{b,-1}$, $Z_{b,-2}$ are Cartier dual to each other under the principal polarization on $\mathbb{X}_{b}$.
@@
    $\delta_{b}$ is an $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$-linear isomorphism
@@
    $Z^{p}$ is an $\mathcal{O}_{B}$ -stable split, symplectic and admissible filtration
@@
There is an action of $J_{b}(\mathbb{Q}_{p})\times\mathbf{G}(\mathbb{A}^{p}_{f})$ on Igusa cusp labels. For the action of $\mathbf{G}(\mathbb{A}^{p}_{f})$ this works exactly as for the Shimura variety (see discussion after Definition 2.3) since we observe that in this case there is also a bijection between tuples $(Z^{p},X,Y,\phi,\varphi_{0},\varphi_{-2})$ and rational versions $(Z^{p}\otimes\mathbb{Q},X,Y,\phi,\varphi_{0}\otimes\mathbb{Q},\varphi_{-2}\otimes\mathbb{Q})$. For the action of $J_{b}(\mathbb{Q}_{p})$, we first observe that by Dieudonné theory we may replace all $p$-divisible groups with their Dieudonné modules, ie. we ask for a filtration
$Z_{b}:0\subset\mathbb{D}(Z_{b,-2})\subset\mathbb{D}(Z_{b,-1})\subset\mathbb{D}(\mathbb{X}_{b}),$
by Dieudonné submodules with $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$-action such that the quotients are also free $W(\overline{\mathbb{F}}_{p})$-modules with $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$-action and isomorphisms
$\delta_{b}:\mathbb{D}(\mathrm{Gr}^{Z_{b}})\simeq\mathbb{D}(\mathbb{X}_{b})\qquad\text{and}\qquad\tilde{\varphi_{0}}:Y\otimes_{\mathbb{Z}}\mathbb{D}(\mathbb{Q}_{p}/\mathbb{Z}_{p})\simeq\mathbb{D}(\mathrm{Gr}^{Z_{b}}_{0}).$
We can consider rational versions obtained by inverting $p$, i.e.
$Z_{b}\left[\frac{1}{p}\right]:0\subset\mathbb{D}(Z_{b,-2})\left[\frac{1}{p}\right]\subset\mathbb{D}(Z_{b,-1})\left[\frac{1}{p}\right]\subset\mathbb{D}(\mathbb{X}_{b})\left[\frac{1}{p}\right],$
and isomorphisms $\delta_{b}\left[\frac{1}{p}\right],\tilde{\varphi_{0}}\left[\frac{1}{p}\right]$, and we see that to obtain a bijection we need to additionally require that $\delta_{b}\left[\frac{1}{p}\right],\tilde{\varphi}_{0}\left[\frac{1}{p}\right]$ restricts to an isomorphism of Dieudonné modules after restricting to the lattice $\mathbb{D}(\mathbb{X}_{b})$. We now see that $J_{b}(\mathbb{Q}_{p})$ clearly acts on rational data with this extra condition.
If $H\subset J_{b}(\mathbb{Q}_{p})\times\mathbf{G}(\mathbb{A}^{p}_{f})$ is a compact open subgroup then an Igusa cusp label at level $H$ is a $H$-orbit of Igusa cusp labels. For a general closed subgroup $H^{\prime}\subset J_{b}(\mathbb{Q}_{p})\times\mathbf{G}(\mathbb{A}^{p}_{f})$, an Igusa cusp label at level $H^{\prime}$ is a compatible family of Igusa cusp labels at level $H$ for all $H^{\prime}\supset H$.
Additionally, we have the following relationship between Igusa cusp labels and cusp labels for the Shimura variety. Given a cusp label $(Z,\Phi)$, we say that an Igusa cusp label $(Z_{b},Z^{p},X,Y,\phi,\varphi_{0},\varphi_{-2},\tilde{\varphi}_{0},\delta_{b})$ lives over $(Z,\Phi)$ if the lattices $X,Y$ are the same, and the base change of $Z$ to $\hat{\mathbb{Z}}^{p}$ recovers the filtration $Z^{p}$ and the isomorphisms $\varphi_{0},\varphi_{-2}$.
Remark 2.8.
The definition in [54, Definition 3.2.15] of an Igusa cusp label is different from ours in that the definition in loc. cit. involves (using our terminology) first fixing a cusp label of level $K=K(N)$, and then considering Igusa cusp labels of level $\{1\}\times K^{p}$ living over this cusp label. The definition of cusp labels in loc. cit. is adapted from Lan’s definition, c.f. the discussion in 2.4. In particular, the Igusa cusp labels in [54] contains extra data of lattices and isomorphisms in $L\otimes\mathbb{Z}_{p}$. This slight variation does not affect any of the results quoted here, since we will always be in a situation where we fix a cusp label $Z$ and are looking at Igusa cusp labels $\tilde{Z}$ living over $Z$.
@@ 2.2.4. Boundary components
We now explain how we can understand the boundary $\partial\mathrm{Ig}^{b,\mathrm{tor}}$ according to Igusa cusp labels following [8] and [54]. Note that since $K=K(N)$ admits a decomposition as $K_{p}K^{p}$, where $K^{p}\subseteq\mathbf{G}(\mathbb{A}_{f}^{p})$, we henceforth use $K^{p}(N)$ to denote this prime-to-$p$ level.
Recall that for a cusp label $Z=(Z,\Phi)$, we have a locally closed boundary stratum $\mathscr{C}_{\mathbb{X}_{b},Z}\subset\mathscr{C}_{\mathbb{X_{b}}}^{\mathrm{tor}}$, and we denote by $\mathrm{Ig}^{b,\mathrm{tor}}_{Z}$ the preimage of $\mathscr{C}_{\mathbb{X}_{b},Z}^{\mathrm{perf}}$ under the map $f:\mathrm{Ig}^{b,\mathrm{tor}}\rightarrow\mathscr{C}_{\mathbb{X_{b}}}^{\mathrm{tor},\mathrm{perf}}$. We want to understand the decomposition of $\mathrm{Ig}^{b,\mathrm{tor}}_{Z}$ according to Igusa cusp labels.
We first recall the following definition in [54] of the perfect boundary components corresponding to the Igusa cusp labels. Set $\mathscr{C}_{C_{Z}}:=\mathscr{C}_{\mathbb{X}_{b},Z}\times_{\mathscr{S}_{Z}}C_{Z}$, and recall that over $\mathscr{C}_{C_{Z}}$ there is a universal Raynaud extension
$0\rightarrow T\rightarrow\mathcal{G}\rightarrow\mathcal{B}\rightarrow 0,$
where $\mathcal{B}$ is an abelian scheme. Similar to above, $\mathcal{B}[p^{\infty}]$ we let the multiplicative part be denoted by $\mathcal{B}[p^{\infty}]^{\mu}$ and the connected part be denoted by $\mathcal{B}[p^{\infty}]^{(0,1)}$.
Definition 2.9.
[54, Definition 3.2.34] Let $\tilde{Z}$ be an Igusa cusp label of level $\{1\}\times K^{p}(N)$ living over a cusp label $Z$ of level $K(N)$. Let $\mathrm{Ig}_{\tilde{Z},C_{Z}}^{b}$ be the functor on perfect schemes $\mathscr{T}$ over $\mathscr{C}_{C_{Z}}$ parametrizing:
  1. (1)
    An isomorphism $\rho:\mathcal{B}[p^{\infty}]^{\mu}\times_{\mathscr{C}_{C_{Z}}}\mathscr{T}\xrightarrow{\sim}\bigl(\mathrm{Gr}^{Z_{b}}_{-1}\bigr)^{\mu}\times_{\overline{\mathbb{F}}_{p}}\mathscr{T}$ compatible with the $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$-actions;
  2. (2)
    A splitting $\delta:\;\mathcal{B}[p^{\infty}]^{\mu}\times_{\mathscr{C}_{C_{Z}}}\mathscr{T}\longrightarrow\mathcal{G}[p^{\infty}]^{\mu}\times_{\mathscr{C}_{C_{Z}}}\mathscr{T}$ compatible with the $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$-actions;
  3. (3)
    An $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$ -linear isomorphism $\rho^{(0,1)}:\mathcal{B}[p^{\infty}]^{(0,1)}\times_{\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}}\mathscr{T}\xrightarrow{\sim}\mathbb{X}^{(0,1)}\times_{\bar{\mathbb{F}}_{p}}{\mathscr{T}}$ which commutes with the polarizations up to the given element of $\mathbb{Z}_{p}^{\times}(\mathscr{T})$.
This functor is representable by a perfect scheme, as implicitly seen from [54]. More precisely, we can first consider a level $p^{m}$-variant $\mathrm{Ig}_{m,\tilde{Z}_{m},C_{Z}}$ as in [54, Definition 3.2.20] for each $m\in\mathbb{N}$ (this is denoted $\mathrm{Ig}^{\mathbb{X}_{b}}_{\tilde{Z}^{m}_{n},C_{\Phi_{n}},\delta_{n}}$ in loc. cit.). Using [8, Theorem 2.2.4] we see that each of these functors is representable by a scheme over $\mathscr{C}_{Z}$, where we further note that in light of (1) the data in (2) of the splitting is equivalent to asking for an isomorphism
$T[p^{\infty}]\oplus\mathcal{B}^{\mu}[p^{\infty}]\simeq\mathcal{G}^{\mu}[p^{\infty}]$
compatible with $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$ actions. Now, taking the perfection of the inverse limit over $m$ recovers the functor defined above, and hence $\mathrm{Ig}_{\tilde{Z},C_{Z}}^{b}$ [^3] This is denoted by $\mathfrak{Ig}^{\mathbb{X}_{b}}_{Z_{n},C_{\Phi_{n}},\delta_{n}}$ in [54], but in this paper we reserve the fraktur font for formal schemes. is also representable by a perfect scheme.
We have a decomposition of $\mathrm{Ig}_{Z}^{b,\mathrm{tor}}$ into locally closed boundary strata indexed by Igusa cusp labels $\tilde{Z}$ at level $K^{p}(N)$ living over a fixed cusp label $Z$. These boundary components are related to $\mathrm{Ig}_{\tilde{Z},C_{Z}}^{b}$ in the following way: Denote by $\widehat{\mathrm{Ig}}^{b,\mathrm{tor}}_{Z}$ the completion of $\mathrm{Ig}^{b,\mathrm{tor}}$ along the locally closed perfect subscheme $\mathrm{Ig}^{b,\mathrm{tor}}_{Z}$. We have the following:
Theorem 2.10.
[54, Theorem 3.2.36]
  1. (1)
    $\widehat{\mathrm{Ig}}^{b,\mathrm{tor}}_{Z}$ admits a decomposition into open and closed formal subschemes
    $\widehat{\mathrm{Ig}}^{b,\mathrm{tor}}_{Z}=\bigsqcup_{\tilde{Z}}\widehat{\mathrm{Ig}}^{b,\mathrm{tor}}_{\tilde{Z}},$
@@
    where $\tilde{Z}$ runs over all Igusa cusp labels of level $\{1\}\times K^{p}(N)$ living over $Z$.
  2. (2)
    Each $\widehat{\mathrm{Ig}}^{b,\mathrm{tor}}_{\tilde{Z}}$ admits the following description: We have a commutative diagram with both squares Cartesian:
    (4) $\mathrm{Ig}^{b}_{\tilde{Z},\Xi}$$\Xi_{Z}^{\mathrm{perf}}$$\mathrm{Ig}^{b}_{\tilde{Z},\Xi,\Sigma_{Z}}$$\Xi_{Z,\Sigma_{Z}}^{\mathrm{perf}}$$\mathrm{Ig}^{b}_{\tilde{Z},C_{Z}}$$C_{Z}^{\mathrm{perf}},$}
    where the composition of the vertical maps are torsors for the torus $E_{\Phi}^{\mathrm{perf}}$, and the inclusion maps in the top square define a relative toroidal embedding.
[…]
    $\widehat{\mathrm{Ig}}^{b,\mathrm{tor}}_{\tilde{Z}}\simeq\widehat{\mathrm{Ig}}^{b}_{\tilde{Z},\Xi,\Sigma_{Z}}/\Gamma_{\tilde{Z}}$,
    Here, we define $\Gamma_{\tilde{Z}}$ to be the subgroup of $\Gamma_{Z}$ consisting isomorphisms $(\gamma_{X},\gamma_{Y})$ such that $\tilde{\varphi}_{0}=\gamma_{Y}\circ\tilde{\varphi}_{0}$.
For later use, we give also an alternate moduli description of $\mathrm{Ig}^{b}_{\tilde{Z},\Xi}$.
Proposition 2.11.
We have the following:
  1. (1)
    1. (a)
      Over $\mathrm{Ig}^{b}_{\tilde{Z},C_{Z}}$, we have universal Raynaud extensions
@@
      and its dual
@@
      equivalently maps $c:X\rightarrow\mathcal{B}$ and $c^{\vee}:Y\rightarrow\mathcal{B}$.
    2. (b)
      Over $\mathrm{Ig}^{b}_{\tilde{Z},C_{Z}}$, there exists a universal $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$-equivariant embedding $\rho:\mathcal{G}[p^{\infty}]\hookrightarrow\mathbb{X}_{b}\times_{\overline{\mathbb{F}}_{p}}\mathrm{Ig}^{b}_{\tilde{Z},C_{Z}}$ which is compatible with polarizations in the following sense: The induced filtration
      $Z_{\mathbb{X}_{b}}:T[p^{\infty}]\subset\mathcal{G}[p^{\infty}]\subset\mathbb{X}_{b}\times_{\overline{\mathbb{F}}_{p}}\mathrm{Ig}^{b}_{\tilde{Z},C_{Z}}$
      is the base change of $Z_{b}$, and the induced isomorphism
      $\mathrm{Gr}^{Z_{b}}_{-1}\simeq\mathcal{G}[p^{\infty}]/T[p^{\infty}]=\mathcal{B}[p^{\infty}]$
      is compatible with the polarizations
    3. (c)
      There is a lift of $c^{\vee}:Y\rightarrow\mathcal{B}$ to $\tilde{c}^{\vee}:Y[\frac{1}{p}]\rightarrow\mathcal{B}$ corresponding to a symplectic splitting of the filtration $Z_{\mathbb{X}_{b}}$.